solve each of the systme by using either addition or substitution. If a unique solution does not exist, state whethere the system is dependent or inconsistent:
10x+2y=7
y=-5x+3

Answer by elima(1433)

10x+2y=7
y=-5x+3
==================
Lets go ahead and use substitution since we already have what y equals, so we will substitute y=-5x+3 in the other equation for y;
10x+2(-5x+3)=7
10x-10x+6=7
6=7
Since both variables are eliminated, the system is inconsistent
:)